# Functional Clustering for Longitudinal Associations between Social Determinants of Health and Stroke Mortality in the US

### v2

@@ 2.4.1 General procedures of REM algorithm
$\bm{\beta}^{(m)}=\text{argmax}_{\bm{\beta}}{Q}^{\text{pen}}\left(\left(\bm{\beta},(\bm{\sigma}^{2})^{(m-1)},\bm{\pi}^{(m-1)}\right)\mid\bm{\Phi}^{(m-1)};\lambda,\rho,r\right).$ (10)
@@
$\pi_{k}^{(m)}=\frac{\sum_{n=1}^{N}\omega_{ik}^{(m)}}{N},\quad k=1,\ldots,K,$ (11)
@@
$(\sigma_{k}^{2})^{(m)}=\frac{\sum_{i=1}^{N}\omega_{ik}^{(m)}\sum_{s=1}^{S_{i}}\left\{Y_{i}(t_{is})-\bm{X}_{i}(t_{is})^{T}\bm{\beta}_{\cdot k}^{(m)}(t_{is})\right\}^{2}}{\sum_{i=1}^{N}S_{i}\omega_{ik}^{(m)}},\quad k=1,\ldots,K,$ (12)
@@
where $\omega_{ik}^{(m)}$ is defined in (7). Note that the main efforts for the above procedures lie in the optimization in (10). We demonstrate this process in detail in Sections 2.4.2.
2.4.2 Optimization of FGSCAD-Net Regularization
For the functional parameter $\bm{\beta}(t)$, we parameterize its $(j,k)$th element, which is $\beta_{jk}(t)$, by the cubic spline basis functions $\psi_{1}(t),\ldots,\psi_{L}(t)$ with equally spaced knots on $\mathcal{T}$. We assume that
$\beta_{jk}(t)=\bm{b}_{jk}^{T}\bm{\Psi}(t),$ (13)
@@ 2.4.2 Optimization of FGSCAD-Net Regularization
where $\bm{b}_{jk}=(b_{jk1},\ldots,b_{jkL})^{T}$ and $\bm{\Psi}(t)=(\psi_{1}(t),\ldots,\psi_{L}(t))^{T}$. We then maximize (10) by substituting (13) into (9). We use $\bm{Q}^{T}\bm{Q}$ to denote the Cholesky decomposition of the non-negative definite matrix
$\int_{\mathcal{T}}\bm{\Psi}(t)\{\bm{\Psi}(t)\}^{T}+r\bm{\Psi}^{{}^{\prime\prime}}(t)\{\bm{\Psi}^{{}^{\prime\prime}}(t)\}^{T}\mathrm{d}t,$
@@
where $\bm{Q}$ is an upper triangular matrix. With this, we define $\bm{\alpha}_{j\cdot}=\left(\bm{\alpha}_{j1}^{T},\ldots,\bm{\alpha}_{jk}^{T}\right)^{T}$ with $\bm{\alpha}_{jk}=\bm{Q}\bm{b}_{jk}$ and rewrite $\|\bm{\beta}_{j\cdot}\|_{r}$ as
@@
$\|\bm{\beta}_{j\cdot}\|_{r}$ $=\sqrt{\|\bm{\beta}_{j\cdot}\|^{2}+r\|(\bm{\beta}_{j\cdot})^{{}^{\prime\prime}}\|^{2}}$
$=\sqrt{\sum_{k=1}^{K}\int\bm{b}_{jk}^{T}\bm{\Psi}(t)\{\bm{\Psi}(t)\}^{T}\bm{b}_{jk}+r\bm{b}_{jk}^{T}\bm{\Psi}^{{}^{\prime\prime}}(t)\{\bm{\Psi}^{{}^{\prime\prime}}(t)\}^{T}\bm{b}_{jk}\mathrm{d}t}$
$=\sqrt{\sum_{k=1}^{K}(\bm{Q}\bm{b}_{jk})^{T}\bm{Q}\bm{b}_{jk}}=\|\bm{\alpha}_{j\cdot}\|,$
where we abuse the notation |||||$\cdot$||| | ⋅ | | to denote the Euclidean norm of a vector. We further denote $\bm{\alpha}_{\cdot k}=\left(\bm{\alpha}_{1k}^{T},\ldots,\bm{\alpha}_{pk}^{T}\right)^{T}$, $h_{ij}(t)=\left(X_{ij}(t)\psi_{1}(t),\ldots,X_{ij}(t)\psi_{L}(t)\right)^{T}\bm{Q}^{-1}$ and $\bm{H}_{i}(t)=(h_{i1}(t)),\ldots,h_{ip}(t))^{T}$. Then we can transform the optimization (10) into the standard group SCAD-$L_{2}$ optimization problem (Zeng and Xie, 2014) as follows
$(\widetilde{\bm{\alpha}}_{\cdot 1}^{(m)},\ldots,\widetilde{\bm{\alpha}}_{\cdot K}^{(m)})=\text{argmax}_{\bm{\alpha}_{\cdot 1},\ldots,\bm{\alpha}_{\cdot K}}Q(\bm{\Phi}\mid\bm{\Phi}^{(m-1)})-\sum_{j=1}^{p}\left\{\bm{P}_{\text{SCAD}}\left(\|\bm{\alpha}_{j\cdot}\|;\lambda\right)+\rho\|\bm{\alpha}_{j\cdot}\|^{2}\right\},$ (14)
where $Q(\bm{\Phi}\mid\bm{\Phi}^{(m-1)})$ can be expressed in terms of $\{\bm{\alpha}_{\cdot k};k=1,\ldots,K\}$ by
$Q(\bm{\Phi}\mid\bm{\Phi}^{(m-1)})=\sum_{i=1}^{n}\sum_{k=1}^{K}\omega_{ik}^{(m)}\sum_{s=1}^{S_{i}}\left[\text{log}\pi_{k}^{(m-1)}+\text{log}\left\{f\left(Y_{i}(t_{is}),\{\bm{H}_{i}(t_{is})\}^{T}\bm{\alpha}_{\cdot k},(\sigma_{k}^{2})^{(m-1)}\right)\right\}\right].$
@@
This optimization problem can be efficiently solved by the group coordinate descent algorithm (Breheny and Huang, 2015).
Since the estimators involved with $L_{2}$ terms are generally biased (Zou and Zhang, 2009; Zeng and Xie, 2014), we conduct a bias correction for $\widetilde{\bm{\alpha}}_{\cdot k}^{(m)}$. In detail, we use the following bias correction formula to obtain the final estimation of $\bm{\alpha}_{\cdot k}^{(m)}$ (Zeng and Xie, 2014)
${\bm{\alpha}}_{\cdot k}^{(m)}=\left\{1+2(1-\rho)\lambda\right\}{\widetilde{\bm{\alpha}}}_{\cdot k}^{(m)}.$ (15)
Once obtaining $\bm{\alpha}_{\cdot k}^{(m)}$, we compute $\bm{\beta}_{\cdot k}^{(m)}(t)$ by
$\bm{\beta}_{\cdot k}^{(m)}(t)=\bm{Q}^{-1}\bm{\alpha}_{\cdot k}^{(m)}\bm{\Psi}(t).$ (16)
2.4.3 Tuning Strategy
@@
The REM algorithm requires tuning of three hyperparameters $\lambda$, $\rho$, and $r$. The traditional strategy for choosing $\lambda$, $\rho$, and $r$ needs to run the entire algorithm for all candidate combinations of $(\lambda,\rho,r)$. This is computationally inefficient since there are numerous choices of candidate combinations that need to be examined. Inspired by the path-fitting algorithm (Breheny and Huang, 2015) for a fast tunning of $\lambda$, we modify the traditional tuning strategy for the REM algorithm. Instead of running a complete REM for each candidate of $\lambda$, we propose to tune $\lambda$ in each M-step of the REM algorithm; refer Part B.2 in Supplementary Material for the implementation details. By this approach, we can employ a path-fitting algorithm to select $\lambda$ for accelerating the tuning process (Huang et al., 2009; Li et al., 2016; Cai et al., 2019).
For the hyperparameters $\rho$ and $r$, we run an entire REM for each of their candidate combinations, nested with the aforementioned strategy for tuning $\lambda$. We select $\rho$ and $r$ from the optional values based on the AIC
$\text{AIC}(\rho,r)=-2l(\widehat{\bm{\Phi}})+2\text{df},$ (17)
where $\widehat{\bm{\Phi}}$ is the converged parameters of the REM algorithm given the current choices of $\rho$ and $r$, and df is the degree of freedom determined as in Breheny and Huang (2015).
In addition, we need to provide the group number $K$ for the implementation of the REM algorithm. To determine $K$, we similarly treat it as a tuning parameter, and proceed with its selection by minimizing the BIC
$\text{BIC}(K)=-2l(\bm{\Phi}_{K})+\text{df}(\lambda)\cdot\text{log}(\sum_{i=1}^{n}S_{i}),$ (18)
where $\bm{\Phi}_{K}$ is the converged parameters when the number of clusters is $K$, with $\rho$ and $r$ selected based on (17). The complete REM algorithm with the tuning processes are summarized in Algorithm 1 in Part B.1 of Supplementary Material.
@@ 3 Simulation
In this section, we conduct numerical simulations to assess the performances of the proposed method in Section 2, in comparison to other competing methods across three aspects: clustering, variable selection, and parameter estimation. To begin with, we generate the functional covariates $\bm{X}_{i}(\cdot)$, $i=1,\ldots,n$, as
$\bm{X}_{i}(t)=\sum_{l=1}^{4}\bm{\theta}_{il}\psi_{l}(t),\ \forall t\in[0,1],$
where $\psi_{1}(\cdot),\ldots,\psi_{4}(\cdot)$ are the first four nonconstant Fourier basis functions, and $\bm{\theta}_{il}\in\mathbb{R}^{p}$, for each $i$ and $l$, is a random vector sampled from a mean-zero Gaussian distribution with the covariance matrix $\left\{l^{-2}\alpha^{|j-k|}\right\}_{1\leq j,k\leq p}\in\mathbb{R}^{p\times p}$. Here, the parameter $\alpha$ controls the dependence between covariates, with a higher value indicating stronger dependencies.
@@
$\bm{\beta}_{\cdot k}(t)=(f_{1k}(t),f_{1k}(t),f_{2k}(t),f_{2k}(t),f_{3k}(t),f_{3k}(t),0,\ldots,0)^{T},\quad k=1,\ldots,K,\ t\in[0,1],$
where $f_{jk}(t)=f_{jk}^{*}(t)/\big{\|}f_{jk}^{*}\big{\|}_{2}$, and $f_{jk}^{*}(\cdot)$s are given by
$f_{11}^{*}(t)=\text{sin}(\frac{\pi t}{2}+\frac{3}{2}\pi)-t-\frac{1}{2},\quad$ $f_{12}^{*}(t)$ =$=\big{\{}\text{cos}(2\pi t)-1\big{\}}^{2},\quad$ $f_{13}^{*}(t)$ $=-f_{11}^{*}(t)+1,$
$f_{21}^{*}(t)=\text{sin}(2\pi t)-t+0.5,\quad$ $f_{22}^{*}(t)$ $=\text{sin}(\frac{\pi t}{2}+\pi),\quad$ $f_{23}^{*}(t)$ $=-f_{21}^{*}(t)-0.5,$
$f_{31}^{*}(t)=-\text{sin}(\frac{\pi t}{2}+\frac{3\pi}{2})-t-0.5,\quad$ $f_{32}^{*}(t)$ $=-f_{12}^{*}(t),\quad$ $f_{33}^{*}(t)$ $=f_{11}^{*}(t)+t+0.5.$
@@
These functions are presented in Figure 2 in Supplementary Material. Note that the relevant covariates for all clusters are $X_{i1},\ldots,X_{i6}$, each of which makes an equal contribution to $Y_{i}$ since $\|f_{jk}\|_{2}=1$ for $j=1,\ldots,6$ and $k=1,\ldots,K$. We set $K=3$.
@@
Next, we generate the cluster membership $Z_{i}$ from a multinomial distribution as described in Section 2, with $\pi_{k}=1/K,\ k=1,\ldots,K$. Providing $\bm{X}_{i}(t),Z_{i}$, and $\bm{\beta}(t)$, $Y_{i}(t)$ is generated from the model (1) for $t$ contained in $10$ equally spaced knots on $[0,1]$, where the variance of the error term $\sigma_{k}^{2}$ is taken according to the signal-to-noise ratio (SNR)
@@
$\sigma_{k}^{2}=\left\{\frac{\sum_{i=1}^{n}\int_{\mathcal{T}}\sum_{k=1}^{K}I(Z_{i}=k)\left\{\bm{X}_{i}(t)^{T}\bm{\beta}_{\cdot k}(t)\right\}^{2}\mathrm{d}t}{n}\right\}\ \bigg{/}\ \text{SNR}.$
@@
We set the SNR as 12. Based on this setting, we evaluate the method proposed in Section 2, which is abbreviated as FGS-Net due to the use of FGSCAD-Net penalty (5). We compare FGS-Net by other REM methods penalized with the FSCAD-Net penalty (4), and FSCAD penalty (3), abbreviated as FS-Net and FS in the following. Apart from these three methods, we examine other competing methods for clustering longitudinal associations.
@@
  • RP: This method incorporates a roughness penalty (RP) into the REM algorithm, which is given as $\sum_{j=1}^{p}\sum_{k=1}^{K}\lambda\|\beta_{jk}^{{}^{\prime\prime}}\|_{2}^{2}$.
  • VS-RP: This method first utilizes the FGS-Net with $K=1$ to conduct the variable selection (VS), where the clustering is not performed at this stage. After that, we adopt the selected variables to implement the RP method for clustering.
@@
  • LI-MIX: This method fits the data by a finite mixture of linear regression models (LI-MIX), i.e., the functional coefficient $\beta_{jk}(t)$ in (1) is treated as a constant over $t$. To preform this method, we first conduct a variable selection using ordinary linear regression models shrunk with an elastic-net penalty (Zou and Zhang, 2009). After that, we adopt the selected covariates to fit a finite mixture of linear regression models for the clustering (Khalili and Chen, 2007).
@@
The RP is a simplification of the FS-Net with $\rho=0$, which only regularizes the roughness of each $\beta_{jk}$ and does not yield sparsity. VS-RP and LI-MIX are two two-step approaches, which implement the variable selection and clustering orderly. These two-step methods are more simple than the aforementioned FGS-Net, FS-Net, FS, and RP. However, selecting relevant covariates prior to the clustering procedure may raise additional problems, as the clustering performance may be sensitive to the outcome from the variable selection. It’s worth noting that LI-MIX further ignores the time-varying nature of $\beta_{jk}$.
@@
For each scenario with different combinations of $n$, $p$, and $\alpha$, the simulations are repeated 100 times. We adopt random initialization for the FGS-Net, FS-Net, and FS methods, i.e., set $\omega_{ik}^{(0)}=I(Z_{i}^{(0)}=k)$, where $Z_{i}^{(0)}$ is sampled from a multinomial distribution as described in Section 2, with $\pi_{k}=1/K$, $k=1,\ldots,K$. On the other hand, RP, VS-RP, and LI-MIX are initialized with the actual cluster membership $\omega_{ik}^{(0)}=I(Z_{i}=k)$. To alleviate computation burdens, we only use one initialization for each simulation. Moreover, the $K$ in RP, VS-RP, and LI-MIX is fixed to 3, the true number of clusters. The $K$ for FGS-Net, FS-Net, and FS are selected based on (18) in Section 2.4.3.
@@
The performance of clustering, variable selection, and parameter estimation is evaluated based on the following criteria.
@@
  • Clustering accuracy is evaluated using the adjusted Rand Index (ARI, Rand, 1971). The ARI is bounded by ±$\pm 1$ to measures the similarity between the true cluster membership and the estimated cluster membership. A higher ARI represents a better clustering result.
@@
  • Variable selection performance is evaluated using C and IC, where C is the number of zero coefficients that are correctly estimated to zero
    $\text{C}=\sum_{j=7}^{p}\sum_{k=1}^{K}I(\|\hat{\beta}_{jk}\|_{2}=0),$
    where $\hat{\beta}_{jk}$ is the estimate of $\beta_{jk}$. Similarly, IC is the number of nonzero coefficients that are incorrectly estimated to zero
    $\text{IC}=\sum_{j=1}^{6}\sum_{k=1}^{{K}}I(\|\hat{\beta}_{jk}\|_{2}=0).$

### v6

@@ 2.4.1 Procedures of REM algorithm
$\bm{\beta}^{(m)}=\operatorname{argmax}_{\bm{\beta}}{Q}^{\text{pen}}\bigl{(}\bigl{(}\bm{\beta},\bigl{(}\bm{\sigma}^{2}\bigr{)}^{(m-1)},\bm{\pi}^{(m-1)}\bigr{)}|\bm{\Phi}^{(m-1)};\lambda,\rho,r\bigr{)}.$ (10)
@@
$\pi_{k}^{(m)}=\frac{\sum_{n=1}^{N}\omega_{ik}^{(m)}}{N},\quad k=1,\ldots,K$ (11)
@@
$\bigl{(}\sigma_{k}^{2}\bigr{)}^{(m)}=\frac{\sum_{i=1}^{N}\omega_{ik}^{(m)}\sum_{s=1}^{S_{i}}\{Y_{i}(t_{is})-\bm{X}_{i}(t_{is})^{\top}\bm{\beta}_{\cdot k}^{(m)}(t_{is})\}^{2}}{\sum_{i=1}^{N}S_{i}\omega_{ik}^{(m)}},\quad k=1,\ldots,K,$ (12)
where $\omega_{ik}^{(m)}$ is defined in (7). Given (10), (11), and (12), we subsequently update $\bm{\Phi}^{(m)}=(\bm{\beta}^{(m)},(\bm{\sigma}^{2})^{(m)},\bm{\pi}^{(m)})$ until the convergence. The stopping criteria of the REM algorithm are provided in Section B of the Supplementary Material.
@@
Note that the main efforts for the above procedures lie in the optimization in (10). We demonstrate this process in detail in Sections 2.4.2.
2.4.2 Optimization of FGS-Net Regularization
Since we penalize the smoothness of $\bm{\beta}$ in (5) via controlling their second-order derivatives, we adopt a similar strategy of smoothing spline (Perperoglou et al. (2019)) to represent the functional coefficients. Specifically, we parameterize $\beta_{jk}(t)$ by the cubic spline basis functions $\psi_{1}(t),\ldots,\psi_{L}(t)$, with the knots of spline functions given as $\bigcup_{i=1}^{n}\{t_{is};s=1,\ldots,S_{i}\}$. As such, $\beta_{jk}(t)$ is represented by
$\beta_{jk}(t)=\bm{b}_{jk}^{\top}\bm{\Psi}(t),\quad t\in[0,1],$ (13)
@@ 2.4.2 Optimization of FGS-Net Regularization
where $\bm{b}_{jk}=(b_{jk1},\ldots,b_{jkL})^{\top}\in\mathbb{R}^{L}$ and $\bm{\Psi}(t)=(\psi_{1}(t),\ldots,\psi_{L}(t))^{\top}\in\mathbb{R}^{L}$. After that, we maximize (10) by substituting (13) into the Q function (9). We use $\bm{D}^{\top}\bm{D}$ to denote the Cholesky decomposition of the nonnegative definite matrix
$\int_{0}^{1}\bm{\Psi}(t)\bigl{\{}\bm{\Psi}(t)\bigr{\}}^{\top}+r\bm{\Psi}^{\prime\prime}(t)\bigl{\{}\bm{\Psi}^{\prime\prime}(t)\bigr{\}}^{\top}\,\mathrm{d}t,$
@@
where $\bm{D}\in\mathbb{R}^{L\times L}$ is an upper triangular matrix. Given this, define $\bm{\alpha}_{jk}=\bm{D}\bm{b}_{jk}\in\mathbb{R}^{L}$ and $\bm{\alpha}_{j\cdot}=(\bm{\alpha}_{j1}^{\top},\ldots,\bm{\alpha}_{jK}^{\top})^{\top}\in\mathbb{R}^{LK}$. We rewrite $\|\bm{\beta}_{j\cdot}\|_{r}$ as
@@
$\lVert\bm{\beta}_{j\cdot}\rVert_{r}$ $=\sqrt{\lVert\bm{\beta}_{j\cdot}\rVert^{2}+r\bigl{\lVert}(\bm{\beta}_{j\cdot})^{\prime\prime}\bigr{\rVert}^{2}}$
$=\sqrt{\sum_{k=1}^{K}\int_{0}^{1}\bm{b}_{jk}^{\top}\bm{\Psi}(t)\bigl{\{}\bm{\Psi}(t)\bigr{\}}^{\top}\bm{b}_{jk}+r\bm{b}_{jk}^{\top}\bm{\Psi}^{\prime\prime}(t)\bigl{\{}\bm{\Psi}^{\prime\prime}(t)\bigr{\}}^{\top}\bm{b}_{jk}\,\mathrm{d}t}$
$=\sqrt{\sum_{k=1}^{K}(\bm{D}\bm{b}_{jk})^{\top}\bm{D}\bm{b}_{jk}}=\lVert\bm{\alpha}_{j\cdot}\rVert,$
[…]
$\operatorname{argmax}_{\bm{\alpha}}\Biggl{[}L_{m}(\bm{\alpha})-\sum_{j=1}^{p}\bigl{\{}\rho\bm{P}_{\text{SCAD}}\bigl{(}\lVert\bm{\alpha}_{j\cdot}\rVert;\lambda\bigr{)}+(1-\rho)\lambda\lVert\bm{\alpha}_{j\cdot}\rVert^{2}\bigr{\}}\Biggr{]},$ (14)
where
$L_{m}(\bm{\alpha})=\sum_{i=1}^{n}\sum_{k=1}^{K}\omega_{ik}^{(m)}\sum_{s=1}^{S_{i}}\Biggl{[}\log\pi_{k}^{(m-1)}+\log\Biggl{\{}f\Biggl{(}Y_{i}(t_{is}),\sum_{j=1}^{p}\bm{h}^{\top}_{ij}(t_{is})\bm{\alpha}_{jk},\bigl{(}\sigma_{k}^{2}\bigr{)}^{(m-1)}\Biggr{)}\Biggr{\}}\Biggr{]}.$
@@
The above optimization can be efficiently solved by the group coordinate descent algorithm (Breheny and Huang (2015)). Once we obtain the maximizer of $\bm{\alpha}_{jk}$ from (14), denoted by ${\bm{\alpha}}^{(m)}_{jk}$, we can estimate the $(j,k)$th element of ${\bm{\beta}}^{(m)}$ by
$\beta_{jk}^{(m)}(t)=\bigl{(}\bm{D}^{-1}\bm{\alpha}_{jk}^{(m)}\bigr{)}^{\top}\bm{\Psi}(t),\quad t\in[0,1].$ (15)
For each $j$, the penalty in (14) encourages ${\bm{\alpha}}_{jk}^{(m)}$, $k=1,\ldots,K$, to be zero vectors simultaneously. The functional coefficients $\{{\beta}_{jk}^{(m)};k=1,\ldots,K\}$ are smooth functions, and they become zero functions simultaneously as $\lambda$ increases, satisfying the cluster-invariant sparsity.
We similarly establish the REM algorithms for the FS penalty (3) or FS-Net penalty (4) by modifying the penalty function in (14) as
[…]
In these two cases, $\{{\bm{\alpha}}_{jk}^{(m)}$, $k=1,\ldots,K\}$ are not encouraged to be zero vectors simultaneously, only obtaining the cluster-variant sparsity during estimation.
2.5 Tuning Strategies
@@
The REM algorithm requires tuning of three hyperparameters $\lambda$, $\rho$, and $r$. The traditional strategy for choosing $\lambda$, $\rho$, and $r$ needs to run the entire algorithm for all candidate combinations of $(\lambda,\rho,r)$. This is computationally inefficient since there are numerous choices of candidate combinations that need to be examined. Inspired by the path-fitting algorithm (Breheny and Huang (2015)) for a fast tunning of $\lambda$, we modify the traditional tuning strategy for our REM algorithm. Instead of running a complete REM for each candidate of $\lambda$, we propose to tune $\lambda$ in each M-step of the REM algorithm; refer to Part B.1 in the Supplementary Material for the implementation details. Using this approach, we can accelerate the tuning process by employing a path-fitting algorithm to select $\lambda$, similar to the existing literature (Huang et al. (2009); Li et al. (2016); Cai et al. (2019)).
For the hyperparameters $\rho$ and $r$, we run an entire REM for each of their candidate combinations, nested with the aforementioned strategy for tuning $\lambda$. Since $\rho$ and $r$ are imposed to penalize functional smoothness and improve prediction performance, we adopt the Akaike Information Criterion (AIC) for the selection of $\rho$ and $r$, as AIC offers a predictive criterion for model selection (Ding et al. (2018)). Specifically, the AIC is calculated as
$\text{AIC}(\rho,r)=-2l\bigl{(}\widehat{\bm{\Phi}}(\rho,r)\bigr{)}+2\text{df}(\rho,r),$ (16)
where $l(\cdot)$ is defined in (2), $\widehat{\bm{\Phi}}(\rho,r)$ is the converged parameters of the REM algorithm, given the current choices of $\rho$ and $r$, and $\text{df}(\rho,r)$ is the degree of freedom determined by $\rho$ and $r$. In addition, we select the number of clusters $K$ using Bayesian information criterion (BIC), a widely used criterion for model-based clustering procedure (Keribin (2000); James and Sugar (2003); Ding et al. (2018); Liang et al. (2021)). The BIC is given as
$\text{BIC}(K)$ $=-2l(\bm{\Phi}_{K})+\text{df}(K)\cdot\log\Biggl{(}\sum_{i=1}^{n}S_{i}\Biggr{)},$ (17)
where $\bm{\Phi}_{K}$ denotes the converged parameter $\widehat{\bm{\Phi}}(\rho,r)$ with $\rho$ and $r$ selected by (16) and the number of clusters being $K$, and $\text{df}(K)$ is the degree of freedom determined by $K$. For the detailed definitions of $\text{df}(\rho,r)$ and $\text{df}(K)$ in AIC (16) and BIC (17), please refer to Breheny and Huang (2015).
@@ 3 Simulation
In this section, we conduct numerical simulations to assess the performances of the proposed method in Section 2, in comparison to other competing methods across three aspects: clustering, variable selection, and parameter estimation.
3.1 Data Generation
We generate the functional covariates $\bm{X}_{i}(\cdot)$, $i=1,\ldots,n$, by
$\bm{X}_{i}(t)=\sum_{l=1}^{4}\bm{\theta}_{il}\psi_{l}(t),\quad\forall t\in[0,1],$
where $\psi_{1}(\cdot),\ldots,\psi_{4}(\cdot)$ are the first four nonconstant Fourier basis functions, and $\bm{\theta}_{il}\in\mathbb{R}^{p}$, for each $i$ and $l$, is a random vector sampled from a mean-zero Gaussian distribution with the covariance matrix $(l^{-2}\alpha^{|j-k|})_{1\leq j,k\leq p}\in\mathbb{R}^{p\times p}$. The parameter $\alpha$ controls the dependence among covariates, with a higher value indicating stronger dependencies.
@@ 3.1 Data Generation
$\bm{\beta}_{\cdot k}(t)=\bigl{(}f_{1k}(t),f_{1k}(t),f_{2k}(t),f_{2k}(t),f_{3k}(t),f_{3k}(t),0,\ldots,0\bigr{)}^{\top},\quad k=1,\ldots,K,t\in[0,1],$
where $f_{jk}(t)=f_{jk}^{*}(t)/\|f_{jk}^{*}\|_{2}$, and $f_{jk}^{*}(\cdot)$s are given by
$f_{11}^{*}(t)$ $=\sin\Bigl{(}\frac{\pi t}{2}+\frac{3\pi}{2}\Bigr{)}-t-\tfrac{1}{2},\quad$ $f_{12}^{*}(t)$ =$=\big{\{}\cos(2\pi t)-1\big{\}}^{2},$
$f_{13}^{*}(t)$ $=-f_{11}^{*}(t)+1,\quad$ $f_{21}^{*}(t)$ $=\sin(2\pi t)-t+0.5,$
$f_{22}^{*}(t)$ $=\sin\Bigl{(}\frac{\pi t}{2}+\pi\Bigr{)},\quad$ $f_{23}^{*}(t)$ $=-f_{21}^{*}(t)-0.5,$
$f_{31}^{*}(t)$ $=-\sin\Bigl{(}\frac{\pi t}{2}+\frac{3\pi}{2}\Bigr{)}-t-0.5,\quad$ $f_{32}^{*}(t)$ $=-f_{12}^{*}(t),$
$f_{33}^{*}(t)$ $=f_{11}^{*}(t)+t+0.5.$
@@
These functions are presented in Figure 2 in the Supplementary Material. It is worth noting that the relevant covariates for all clusters are $X_{i1},\ldots,X_{i6}$, each of which makes an equal contribution to $Y_{i}$ since $\|f_{jk}\|_{2}=1$ for $j=1,\ldots,6$ and $k=1,\ldots,K$. We set the number of clusters $K=3$.
@@
Next, we generate the cluster membership $Z_{i}$ from a multinomial distribution as described in Section 2, with $\pi_{k}=1/K,k=1,\ldots,K$. Providing $\bm{X}_{i}(t),Z_{i}$, and $\bm{\beta}(t)$, $Y_{i}(t)$ is generated from the model (1) for each $t$, where the $\varepsilon_{ik}(t)$ is constructed as $\tilde{\varepsilon}_{ik}(t)+\tau_{ik}$. For each $i$ and $k$, $\{\tilde{\varepsilon}_{ik}(t);t\in[0,1]\}$ are independent mean-zero Gaussian processes with the covariance function ${\sigma_{k}^{2}}\operatorname{exp}(-|t_{1}-t_{2}|)/2$, and $\tau_{ik}$s are Gaussian measurement noises with variance $\sigma_{k}^{2}/2$. The variance of the error terms $\sigma_{k}^{2}$ is taken according to the signal-to-noise ratio (SNR)
@@
$\sigma_{k}^{2}=\biggl{\{}\frac{\sum_{i=1}^{n}\int_{0}^{1}\sum_{k=1}^{K}I(Z_{i}=k)\{\bm{X}_{i}(t)^{\top}\bm{\beta}_{\cdot k}(t)\}^{2}\,\mathrm{d}t}{n}\biggr{\}}/\text{SNR},$
with SNR being set as 12. Finally, the observed time points $\{t_{is};s=1,\ldots,S_{i}\}$ for functional data are set as 10 equally spaced knots on $[0,1]$ for each $i$, mimicking the setting of our SDOH data.
@@
Based on the above setting, we evaluate the method proposed in Section 2, which is denoted as FGS-Net due to the use of FGS-Net penalty (5). We compare FGS-Net by other REM methods penalized with the FS-Net penalty (4) and FS penalty (3), denoted as FS-Net and FS in what follows. Apart from these three methods, we examine the following competing methods for clustering longitudinal associations:
@@
  • RP: This method incorporates a roughness penalty (RP) into the REM algorithm, which is given as $\sum_{j=1}^{p}\sum_{k=1}^{K}\lambda\|\beta_{jk}^{\prime\prime}\|_{2}^{2}$.
  • VS-RP: This method initially employs the FGS-Net within the finite mixture of functional linear concurrent model (1), setting $K=1$ for variable selection (VS). This method is denoted as FMFLCM(1), encompassing several existing methods (Wang et al. (2008); Goldsmith and Schwartz (2017); Ghosal et al. (2020)) as special cases. Subsequently, the preselected variables are used to implement the RP method for clustering.
@@
  • LI-MIX: This method fits the data by a finite mixture of linear regression models (LI-MIX), that is, the functional coefficient $\beta_{jk}(t)$ in (1) is treated as a constant over $t$. To preform this method, we first conduct a variable selection using conventional linear regression models shrunk with an elastic-net penalty (Zou and Zhang (2009)). After that, we adopt the selected covariates to fit a finite mixture of linear regression models for the clustering (Peel and MacLahlan (2000); Khalili and Chen (2007); Khalili and Lin (2013)).
@@
The RP is a simplification of the FS-Net penalty (4) with $\rho=0$, which only regularizes the roughness of each $\beta_{jk}$ and does not yield sparsity. VS-RP and LI-MIX are two two-step approaches, which implement the variable selection and clustering orderly. These two-step methods are more simple than the aforementioned FGS-Net, FS-Net, FS, and RP. However, selecting relevant covariates prior to the clustering procedure may raise additional problems, as the clustering performance may be sensitive to the outcome from the variable selection. It is worth noting that the two-step method LI-MIX further ignores the time-varying nature of $\beta_{jk}$.
@@
For each scenario with different combinations of $n$, $p$, and $\alpha$, the simulations are repeated 100 times. We adopt random initialization for the FGS-Net, FS-Net, and FS methods by setting $\omega_{ik}^{(0)}=I(Z_{i}^{(0)}=k)$, where $Z_{i}^{(0)}$ is sampled from a multinomial distribution over $\{1,\dots,K\}$ with equal probabilities. On the other hand, RP, VS-RP, and LI-MIX are initialized with the actual cluster membership $\omega_{ik}^{(0)}=I(Z_{i}=k)$. To alleviate computation burdens, we only use one initialization for each simulation. Moreover, the $K$ in RP, VS-RP, and LI-MIX is fixed to 3, the true number of clusters. The $K$ for FGS-Net, FS-Net, and FS are selected based on (17).
@@
The performance of clustering, variable selection, and parameter estimation is evaluated based on the following criteria:
@@
  • Clustering accuracy is evaluated using the adjusted Rand Index (ARI, Rand (1971)). The ARI is bounded by ±$\pm 1$ to measures the similarity between the true cluster membership and the estimated cluster membership. A higher ARI represents a better clustering result.
@@
  • Variable selection performance is evaluated using C and IC, where C is the number of zero coefficients that are correctly estimated to zero $\text{C}=\sum_{j=7}^{p}\sum_{k=1}^{K}I\bigl{(}\lVert\hat{\beta}_{jk}\rVert_{2}=0\bigr{)},$ where $\hat{\beta}_{jk}$ is the estimate of $\beta_{jk}$. Similarly, IC is the number of nonzero coefficients that are incorrectly estimated to zero $\text{IC}=\sum_{j=1}^{6}\sum_{k=1}^{{K}}I\bigl{(}\lVert\hat{\beta}_{jk}\rVert_{2}=0\bigr{)}.$